The path of 2 objects are described by the parametric equations below.

object 1: 

x = t

y = t2 + 2


object 2:

 x= 2t +1

y= t + 4



Find when and where the objects collide.

They collide at (_,_) when t = 

.


1
Expert's answer
2021-03-31T16:27:41-0400

Graph of object 1 (G1)


"x=t\\" .......(1)

"y=t^2+2" .........(2)


Putting 1 in 2 we get,

"\boxed{y=x^2+2}" ..…....(G1)


Now Graph of object 2 (G2)…


"x=2t+1" ..........(3)

"y=t+4" ...........(4)


Putting (3) in(4)..we get,

"\boxed{y={x+7\over 2}}" ............(G2)


Now we find intersection of G1 and G2,





We get,

"\boxed{y=4.25,

X=1.5}"


As point of intersection,


For t we put y coordinates of both objects are equal,


"t^2+2=t+4"


we get,

t=-1,2

But -1 can't be expected because t can't be negative.

So "\boxed{t=2}" is answer.



They collide at (1.5,4.25) when t = 2.